# Parameter-Efficient Fine-Tuning of State Space Models

### v1

@@ D.2 Extension to Deep S4 Models
Since SDT is a special case of SDLoRA, Theorem 10 directly implies the desired statement. ∎
SDLoRA for Mamba.
In the Mamba model, the output mapping vector ${\bm{C}}$ is input-dependent, making it unsuitable for direction modification. Therefore, we focus our channel and hidden state selection solely on ${\bm{A}}$. For any channels or hidden states that are not selected, we set the corresponding elements of ${\bm{A}}$ to minimal values, effectively setting the associated entries in $\overline{{\bm{A}}}$ to zero. For channels and states that are updatable, we update the corresponding entries for ${\bm{A}}$. However, since ${\bm{B}}^{(d)}$ and ${\bm{C}}^{(d)}$ cannot be directly updated, we modify the corresponding weight matrices that compute these vectors. Specifically, for updatable channels, we update the corresponding columns in ${\bm{W}}_{{\bm{B}}}$ and ${\bm{W}}_{{\bm{C}}}$; for updatable states, we adjust the corresponding rows in these weight matrices.
D.3 Experiments on Deep S4 Models
@@
We additionally consider SDT (Selective Dimension Tuning), which is introduced in Sec. D.2, and the results are visualized in Fig. 2. We observe that SDT even outperforms SDLoRA in this synthetic experiments, demonstrating highly promising performance. Unfortunetly, we fail to make it work on pretraned Mamba, and identify it as one of the promising future directions.
Figure 2: Performance comparison between various methods. SDT (Selective Dimension Tuning) is compared to SDLoRA. Unlike SDLoRA, which applies LoRA to linear projection matrices, SDT performs sparse tuning on linear projection matrices by updating only the columns corresponding to channels selected via Alg. 1. Notably, SDT achieves superior performance, matching full fine-tuning results while using only 25% of the parameters, and even surpassing full fine-tuning with more parameters. Extending SDT to real datasets is considered a promising future direction for SDLoRA.
@@
We adhere to the preprocessing steps for CIFAR-10 as outlined by Gu et al. (2022a). The LoRA ranks for linear projection matrices are tuned among {1, 2, 4, 8, 16}, and for the S4 component, ranks are set from {1, 2, 4}. Non-zero states are chosen from {8, 12, 16}, and non-zero channels from {48, 64}. A warmup phase includes 1 epoch with a learning rate of $21e-$ 2. For linear projection matrices, LoRA ranks are explored at {2, 4, 8, 16}, and for the SSM, ranks at {2, 4, 8}. All state dimensions are updated, and channel dimensions considered for updates are {4, 8, 16, 32}.
D.4 Experiments on Pretrained Mamba
@@
Here, we provide more experiment details. Unless otherwise stated, our experiment setting is identical to Sec. C.1. For LoRA, we consider three different LoRA configurations at each layer, involving the following matrices which comprise most of the parameters: $\bm{W}_{\text{out}}$ (output linear projection), $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ (weight matrices for computing input-dependent ${\bm{B}}_{n},{\bm{C}}_{n}$), and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$ (down and up projection matrices of LoRA adapters for computing $\bm{\Delta}$). The three LoRA application methods are: (i) $\bm{W}_{\text{out}}$, $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$; (ii) $\bm{W}_{\text{out}},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ and ${\bm{W}}_{\bm{\Delta},\downarrow}$; and (iii) $\bm{W}_{\text{out}}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$. For SDLoRA, we set the channel freeze ratio at 99% across all scenarios. We select the state freeze ratio $\alpha$ from the set ${75\%,90\%,95\%}$ and apply LoRA exclusively to $\bm{W}_{\text{out}}$ to maintain a comparable number of trainable parameters. Residual connections and bias are frozen in this experiment. For the warmup, we employ 500 data batches to fully train the SSM modules prior to dimension selection, except for the RTE task in GLUE, where we use 250 batches due to its limited dataset size. Note that the parameters are reverted back after the warmup stage.

### v3

@@ D.2 Extension to Deep S4 Models
Since SDT+ is a special case of SDT-P, Theorem 2 directly implies the desired statement. ∎
D.3 Extension to S6
In this section, we extend the discussion of SDT-P and SDT+ to S6, following the same logic. We begin by proving results for SDT+ in the scenario where the target model consists of only a single layer. In doing so, we extend Theorem 3 to apply to deep S6 models by first generalizing Lemma 5 to Lemma 6.
Lemma 6.
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S6 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target one-layer deep S6 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$. Then this can be achieved by selectively fine-tuning at most $\lceil D/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Proof of Lemma 6.
To demonstrate this, we can follow the same proof strategy as in the proof of Lemma 5. In particular, the $t$-th intermediate output token of the $l$-th deep S6 layer of the updated model can be similarly written as
$\begin{split}{\bm{z}}_{l,t}&=\widehat{{\bm{W}}}_{l}\cdot\widehat{\operatorname{S6}}_{l,t}({\bm{z}}_{l-1,1},\ldots,{\bm{z}}_{l-1,t})+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t}\\ &=\widehat{{\bm{W}}}_{l}\cdot\begin{bmatrix}\widehat{\operatorname{S6}}_{l,t}^{(1)}(z_{l-1,1}^{(1)},\ldots,z_{l-1,t}^{(1)})\\ \vdots\\ \widehat{\operatorname{S6}}_{l,t}^{(D)}(z_{l-1,1}^{(D)},\ldots,z_{l-1,t}^{(D)})\\ \end{bmatrix}+\widehat{\bm{\beta}}_{l}+\widehat{{\bm{u}}}_{l}\odot{\bm{z}}_{l-1,t},\end{split}$ (57)
For layers $l<L=D$.
  • (Channel Selection) Select $D^{\prime}\leq D$ ($D^{\prime}=1$ here) important channels for making predictions. For any channel $d$ that is not utilized, rather than directly setting the corresponding ${\bm{C}}^{(d)}$ to zero as in the deep S4 model, we instead set $\bm{\beta}^{(d)}_{\bm{\Delta}}$ to be sufficiently large. According to the computation of SSM parameters described in (58), this ensures that $\overline{{\bm{B}}}_{n}^{(d)}$ is set to zero for all $d\neq l$ in this scenario. This approach is equivalent to setting ${\bm{C}}^{(d)}$ to zero, as both result in the channel producing all zeros.
  • (Hidden State Selection) Within the selected channels, select $H^{\prime}\leq H$ important hidden states. For any hidden state that is not used within a selected channel $d$, the corresponding entries in ${\bm{A}}^{(d)}$ will be set to sufficiently small. To be more specific, we can achieve $\widehat{\operatorname{S6}}_{l,t}^{(l)}(\cdot)=\operatorname{S6}_{\star,t}^{(l)}(\cdot)$ by leveraging the discretized parameters. Lemma 2 provides the conditions for this equality to hold by updating $\overline{{\bm{A}}}$, $\overline{{\bm{B}}}$, and ${\bm{C}}$ for the $l$-th channel, where these parameters are computed as follows:
    $\bm{\Delta}_{n}=\operatorname{softplus}({\bm{W}}_{\bm{\Delta},\downarrow}{\bm{W}}_{\bm{\Delta},\uparrow}\bm{x}_{n}+\bm{\beta}_{\bm{\Delta}}),$ (58)
    $\overline{{\bm{A}}}_{n}^{(d)}=\exp(\Delta_{n}^{(d)}{\bm{A}}^{(d)}),\quad\overline{{\bm{B}}}_{n}^{(d)}=\Delta_{n}^{(d)}\bm{W}_{{\bm{B}}}\bm{x}_{n},\quad{\bm{C}}_{n}=\bm{W}_{{\bm{C}}}\bm{x}_{n},$ (59)
    Therefore, we can achieve $\widehat{\operatorname{S6}}_{l,t}^{(l)}(\cdot)=\operatorname{S6}_{\star,t}^{(l)}(\cdot)$ by only updating the corresponding values or columns of the weight matrices for each channel and dimension.
  • (Residual and Bias Tuning) Regardless of other selections, SDT+ consistently tunes the coefficients of residual connections and biases in linear projections, as these components contain a negligible number of parameters. In this scenario, we let $\widehat{\bm{\beta}}_{l}={\bm{0}},\leavevmode\nobreak\ \widehat{{\bm{u}}}_{l}=\begin{bmatrix}\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{l-1\text{ elements}}}&0&\smash{\underbrace{\begin{matrix}1&\cdots&1\end{matrix}}_{D-l\text{ elements}}}\end{bmatrix}^{\top}.$
This construction yields
For the remaining layers, following the same steps leads to the desired results. ∎
Therefore, we similarly obtain the following two results.
Theorem 4 (Expressive Power of SDT+ on Deep S6 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT+, any deep S6 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S6 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states, and residual connections at each layer.
Theorem 5 (Expressive Power of SDT-P on Deep S6 Models).
Consider a $D$-dimensional input sequence. Assume that the linear layers in the model have linear activation functions. Using SDT-P, any deep S6 model with $H$ hidden states per channel and $L$ layers can be updated to accurately present any target deep S6 model without residual connections, having a reduced hidden state dimension $H^{\star}<H$, and fewer layers $L^{\star}<L$. This can be achieved by selectively fine-tuning at most $\lceil DL^{\star}/L\rceil$ channels, $H^{\star}$ hidden states on SSM modules, applying rank-$\lceil\frac{L}{L^{\star}}\rceil$ updates on linear projection matrices and updating residual connections and biases at each layer, while additionally fully fine-tuning the linear projection matrix of the last layer only.
Combining Theorem 3 and 5 leads to Theorem 1.
D.4 Sparse Dimension Tuning and Pruning (SDT-P)
Algorithm 2 is our extended algorithm, which includes setting dimensions to zero. However, in practical settings, setting channels to zero is not necessary and omitting it reduces number of hyperparameters, as pruning parameters is effectively equivalent to training them to zero.
Algorithm 2 Dimension Selection Algorithm of SDT-P
Input: A small subset of dataset ${\mathcal{D}}$, warmup epochs $E$, number of layers $L$, total channels $D$, total states $H$, state sparsity $\beta_{0}$, channel sparsity $\alpha_{0}$, state update fraction $\beta$, channel update fraction $\alpha$
/* Warmup Epochs */
Perform full update on SSM modules using ${\mathcal{D}}$ for $E$ epochs  /* Categorize dimensions */
for $l=1$ to $L$ do
       /* Set dimensions as zero */ Sort channels based on $\|\overline{{\bm{A}}}^{(d)}\|$   Select final $(1-\beta_{0})D$ channels as zero channels and denote non-zero channels as set ${\mathbb{D}}$for $d\in{\mathbb{D}}$ do
             Sort states based on magnitude of $\bar{A}^{(d)}_{h}$ at each state dimension  Select final $(1-\alpha_{0})H$ states as zero states and denote non-zero states as set ${\mathbb{H}}$
      /* Unfreeze dimensions */ Sort non-zero channels ${\mathbb{D}}$ based on changes of $\|\overline{{\bm{A}}}^{(d)}\|$  Select the top $\beta|{\mathbb{D}}|$ channels as updatable, denoted by ${\mathbb{D}}^{\prime}$for $d\in{\mathbb{D}}^{\prime}$ do
             Sort non-zero state dimensions based on changes of $\|\overline{{\bm{A}}}^{(d)}\|$  Select the top $\alpha|{\mathbb{H}}|$ states as updatable at the $d$-th channel
D.5 Extension to S5
In this part, we extend Lemma 2 and Theorem 1 to two corresponding results for S5. The extension follows the same procedure as the previous proof, so we omit the details here.
Lemma 7 (Minimal Parameter Adjustment for S5 Fine-Tuning).
Assume all hidden dimensions of the target model $f^{\star}$ are non-zero, i.e., all elements of $\operatorname{diag}(\overline{{\bm{A}}}_{\star})\odot\overline{{\bm{B}}}_{\star}^{(d)}\odot{\bm{C}}_{\star}^{(d)}$ are non-zero. To update frozen model $f_{0}$ such that it becomes functionally equivalent to the target model $f^{\star}$, the minimum number of tunable parameters is: $\min_{\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}}}\left\lVert\left[\overline{{\bm{A}}}\right]_{1:H_{\star},1:H_{\star}}-\overline{{\bm{A}}}_{\star}\right\rVert_{0}+\sum_{d=1}^{D}\left(\overbrace{\left\lVert\left[\operatorname{diag}(\overline{{\bm{A}}})\odot\overline{{\bm{B}}}^{(d)}\odot{\bm{C}}^{(d)\top}\right]_{(H_{\star}+1):H}\right\rVert_{0}}^{\text{eliminatingredundant dimensions}}+\overbrace{\left\lVert\left[\overline{{\bm{B}}}^{(d)}\odot{\bm{C}}^{(d)\top}\right]_{1:H_{\star}}-\overline{{\bm{B}}}_{\star}^{(d)}\odot{\bm{C}}_{\star}^{(d)\top}\right\rVert_{0}}^{\text{ aligning remaining dimensions with target model}}\right),$
subject $\textnormal{subject to}\quad(\overline{{\bm{A}}},\overline{{\bm{B}}},{\bm{C}})\in\{({\bm{P}}^{\top}\overline{{\bm{A}}}_{0}{\bm{P}},{\bm{P}}^{\top}\overline{{\bm{B}}}_{0},{\bm{C}}_{0}{\bm{P}}):{\bm{P}}\text{ is a permutation matrix}\}.$ (61)
Theorem 6 (Expressive Power of SDT-P with LoRA on Simplified SSM-based Models).
Assume all layers use linear activations. Let $f_{0}$ be a frozen deep S4 S5, or S6 model with $L$ layers, each containing $H$ hidden states per channel. Let $f^{\star}$ be a smaller target model of the same type (S4, S5 or S6), with no residual connections, $L^{\star}<L$ layers, and $H^{\star}<H$ hidden states per channel. Then, there exists a set of parameter updates to $f_{0}$ satisfying the following conditions such that for any finite-length input sequence ${\bm{X}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{N})$ with ${\bm{x}}_{n}\in{\mathcal{X}}\subset{\mathbb{R}}^{D}$, where ${\mathcal{X}}$ is bounded, the resulting model $f$ satisfies $f({\bm{X}})=f^{\star}({\bm{X}})$:
  1. 1.
    (SDT-P on SSM) In each SSM module, update at most $\lceil DL^{\star}/L\rceil$ channels. Within each updated channel, fine-tune at most $H^{\star}$ hidden states and set the rest to zero.
  2. 2.
    (LoRA on Linear Projections) Apply rank-$\lceil L/L^{\star}\rceil$ updates to each linear projection matrix.
  3. 3.
    (Minimal Additional Updates) Modify only the residual connections, per-layer biases, and the final-layer output projection.
D.6 Memory Usage and Runtime Analysis of SDT
Memory Usage Analysis.
To assess the memory usage of SDT and LoRA, we conducted experiments on four different models, including both SSM and hybrid architectures. For each model and method, a dataset was generated with 2,500 batches of data samples, each batch comprising a random sequence of 1,500 tokens. The simulation was repeated four times, including dataset generation. All experiments were carried out on a single H100 GPU, and the reported metrics represent averages across the four simulations. Consistent with our previous experiments, we used the original hyperparameter settings, ensuring that SDT includes similar trainable parameters than LoRA. The memory usage of LoRA and SDT is presented in Table 16. Our observations indicate that SDT requires less memory than LoRA. This difference can be attributed to the design of the LoRA adapters, which involve matrix multiplication of two low-rank matrices. In contrast, tuning SSM with the same number of parameters does not require any matrix multiplication, resulting in lower memory usage.
Memory Usage (GB) Mamba-130M Mamba-1.4B Jamba-Tiny-319M Jamba-Mini-52B
LoRA 7.753 37.167 7.207 71.986
LoRA & SDT 5.738 26.491 6.605 67.193
Table 16: Memory usage comparison between SDT and LoRA on various models. Bold numbers indicate the lowest memory usage for each model.
To provide a more fine-grained view, we further analyze how sequence length affects peak memory usage for different Mamba model sizes, as shown in Fig. 4. We measure the memory required to process a single training batch with varying context lengths using randomly generated data. Each batch contains four examples, with 90% of tokens used as input and 10% as output (loss is computed only on the output tokens). The experimental settings for both LoRA and SDT follow the setup described in Section 6.2. We evaluate three configurations for each method, matched in parameter budget. In the plot, each line represents the average across the three configurations, and the shaded region for LoRA shows the range (minimum to maximum). SDT shows negligible variance across configurations, so no shading is included. All models are trained for 500 iterations, and results are averaged over these iterations. Experiments were conducted on an NVIDIA H100 80GB GPU.
Figure 4: Peak memory usage during training as a function of context length for different Mamba model sizes. Each line represents the mean across three configurations; shaded regions indicate min–max ranges. SDT is consistently more memory-efficient than LoRA when applying to SSM module.
Runtime Analysis.
We similarly analyze the latency of LoRA and SDT, using the same experimental setup as in Table 16. Fine-tuning with SDT consists of two stages: (1) dimension selection and (2) standard training. In this study, we first compare the runtime of SDT and LoRA during stage 2 (training) and then evaluate the additional runtime introduced by SDT during stage 1 (dimension selection). Our results show that the dimension selection stage adds only marginal runtime overhead, and SDT is more efficient than LoRA in standard training.
Training: When the channels and states have been selected, the training of SDT is faster than LoRA when the same number of trainable parameters are considered.
The runtimes are reported in Table 17. We observe that, despite having similar trainable parameters, SDT is faster than LoRA. We attribute this to the fact that LoRA introduces additional FLOPs due to the extra matrix multiplication operations required for each update (specifically, the multiplication of two low-rank matrices).
Avg. Runtime (Seconds) Mamba-130M Mamba-1.4B Jamba-Tiny-319M Jamba-Mini-52B
LoRA 410.0 $\pm$ 80.0 2060.0 $\pm$ 135.0 352.5 $\pm$ 107.5 3427.5 $\pm$ 185.0
LoRA & SDT 330.0 $\pm$ 77.5 1697.5 $\pm$ 87.5 257.5 $\pm$ 72.5 3065.0 $\pm$ 232.5
Table 17: Runtime comparison of SDT and LoRA during stage 2 (training).
Dimension Selection: For dimension selection, our method first performs Initial Subset Training, and then selects the dimensions based on the magnitude of parameter changes across different dimensions.
  1. 1.
    Initial Subset Training: We update the model by going through only a subset of the dataset (e.g., 3% of batches in DART experiments), which is sufficient in practice.
  2. 2.
    Magnitude-Based Dimension Selection: After the subset training, we select dimensions based on the magnitude of parameter changes observed.
In this experiment, we simulate a real scenario using a dataset with 2,500 batches, considering a small subset containing 125 batches (5% of the full dataset). We repeat the experiments 80 times, and the reported numbers are averaged across these simulations. Table 18 demonstrates that the dimension selection stage adds only negligible runtime.
Avg. Runtime (Seconds) Mamba-130M Mamba-1.4B Jamba-Tiny-319M Jamba-Mini-52B
Initial Subset Training 16.250 $\pm$ 3.880 85.250 $\pm$ 5.130 15.750 $\pm$ 1.000 163.630 $\pm$ 10.120
Magnitude-Based Dimension Selection 0.280 $\pm$ 0.000 0.520 $\pm$ 0.120 0.090 $\pm$ 0.000 0.240 $\pm$ 0.040
Total Time 16.530 $\pm$ 3.880 85.770 $\pm$ 5.250 15.840 $\pm$ 1.000 163.870 $\pm$ 10.160
Proportion of Training 1 Epoch 0.050$\times$ 0.051$\times$ 0.062$\times$ 0.053$\times$
Proportion of Training 5 Epoch 0.010$\times$ 0.010$\times$ 0.012$\times$ 0.011$\times$
Table 18: Runtime comparison of SDT and LoRA during stage 1 (dimension selection).
We further examine how runtime varies with sequence length, using the same experimental setup as in the memory analysis (Fig. 4). Our results in Fig. 5 show that SDT consistently outperforms LoRA in training speed when applied to SSM modules.
Figure 5: Average training time per batch across different sequence lengths and Mamba model sizes. Each line represents the mean across three configurations; shaded regions indicate min–max ranges. SDT is consistently faster than LoRA when applying to SSM module.
Appendix E Expanded Sec. 6: Evaluation of SDT
E.1 Experiments on Deep S4 Models
@@
In addition, we compare the convergence speed of LoRA and SDT in terms of training loss for sequence lengths in $\{100,500,1000\}$. We plot the MSE of both methods against wall-clock time. As shown in Fig. 6, SDT consistently converges to a lower loss faster than LoRA across all tested sequence lengths.
Figure 6: Comparison of SDT and LoRA for tuning S4 in deep S4 models, where LoRA is applied to linear projection matrices. Results are shown across varying sequence lengths under different time budgets in synthetic experiments.
@@
Previous work (Dinh et al., 2022) demonstrates that large language models can be fine-tuned for image classification tasks. In this study, we investigate the adaptation of SSMs for computer vision, focusing on experiments conducted with the CIFAR-10 dataset (Krizhevsky et al., 2009). We employ an eight-layer deep S4 model with a hidden state dimension of 16 and a model dimension of 64. Since pretrained deep S4 models are not available, we simulate a pretrained scenario by fully updating the model for 50 epochs first, then subsequently evaluating the PEFT methods over an additional 5 epochs. We adhere to the preprocessing steps for CIFAR-10 as outlined by Gu et al. (2022a). The LoRA ranks for linear projection matrices are tuned among {1, 2, 4, 8, 16}, and for the S4 component, ranks are set from {1, 2, 4}. Non-zero states are chosen from {8, 12, 16}, and non-zero channels from {48, 64}. The warmup phase includes 1 epoch with a learning rate of $21e-$ 2. For linear projection matrices, LoRA ranks are explored at {2, 4, 8, 16}, and for the SSM, ranks at {2, 4, 8}. All state dimensions are updated, and channel dimensions considered for updates are {4, 8, 16, 32}. The results, as reported in Table 19, indicate that SDT outperforms LoRA with fewer trainable parameters.
Method # Params (%) Accuracy
Frozen 0.00 73.9
LoRA (Proj) 16.00 77.6
LoRA (S4+Proj) 15.52 77.6
LoRA & SDT 11.17 78.0
Full Fine-Tuning 100.00 77.6
Table 19: Accuracy comparison between SDT and LoRA on deep S4 models for CIFAR-10 (Krizhevsky et al., 2009).
E.2 Experiments on Mamba-II, Jamba, and LoRA$+$
Additional Experimental Details.
@@ E.2 Experiments on Mamba-II, Jamba, and LoRA$+$
In this paragraph, we provide further experimental details. Unless otherwise stated, our experiment setting is identical to Sec. C.1. For LoRA, we consider three different LoRA configurations at each layer, targeting the primary parameters of Mamba. Specifically, we focus on the following matrices: $\bm{W}_{\text{out}}$ (output linear projection), $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ (weight matrices for computing input-dependent ${\bm{B}}_{n},{\bm{C}}_{n}$), and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$ (down and up projection matrices of LoRA adapters for computing $\bm{\Delta}$). The three LoRA application methods are: (i) $\bm{W}_{\text{out}}$, $\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$, and ${\bm{W}}_{\bm{\Delta},\downarrow},{\bm{W}}_{\bm{\Delta},\uparrow}$; (ii) $\bm{W}_{\text{out}},\bm{W}_{{\bm{B}}},\bm{W}_{{\bm{C}}}$ and ${\bm{W}}_{\bm{\Delta},\downarrow}$; and (iii) $\bm{W}_{\text{out}}$ and ${\bm{W}}_{\bm{\Delta},\uparrow}$. For SDT, we set the channel freeze ratio at 99% across all scenarios. We select the state freeze ratio $\alpha$ from the set $\{75\%,90\%,95\%\}$ and apply LoRA exclusively to $\bm{W}_{\text{out}}$ to maintain a comparable number of trainable parameters. Residual connections and biases are frozen in this experiment. For the warmup, we employ 500 data batches to fully train the SSM modules prior to dimension selection, except for the RTE task in GLUE, where we use 250 batches due to its limited dataset size. Note that the parameters are reverted back after the warmup stage.
Additional Results on Mamba-II.
For Mamba-II, applying SDT is not straightforward because Mamba-II further constrains $A$ such that all (non-zero) entries must have the same value. Therefore, our original dimension selection approach cannot be directly applied here. We consider a naive extension of SDT by selecting dimensions in the projection matrices for input mapping vector $B$ and the projection matrices for output mapping vector $C$ using their respective magnitude, and fine-tune the selected dimensions and all elements of state transition matrix $A$.
Tables 20 and 21 compare the performance on Mamba-II. The results demonstrate that SDT consistently outperforms LoRA on Mamba-II models.
Model Mamba-II-130M Mamba-II-1.3B
Dataset Params (%) DART Params (%) SAMSum Spider
Metric ($\uparrow$) METEOR BLEU R1 R2 RL Acc.
LoRA 0.3354 68.71 48.09 0.1614 49.73 26.14 41.53 72.36
LoRA & SDT 0.3393 70.60 48.93 0.1767 50.72 27.21 42.54 84.15
Table 20: Performance comparison between SDT and LoRA on Mamba-II-130M and Mamba-II-1.3B. Bold numbers indicate the best performance for each task.
Model Mamba-II-130M
Dataset Params (%) GLUE
Accuracy ($\uparrow$) RTE MRPC SST2 QNLI QQP MNLI
LoRA 0.3354 63.4 80.9 89.1 85.3 87.1 78.6
LoRA & SDT 0.3393 64.3 82.3 94.1 87.0 88.3 81.1
Table 21: Performance comparison between SDT and LoRA on the GLUE (Wang et al., 2019) benchmark using Mamba-II-130M. Bold numbers indicate the best performance for each task.
Additional Results on Jamba.
Table 22 shows results for SDT and LoRA on additional datasets. Even though the performance improvement is smaller, our method outperforms pure LoRA in most cases. Mamba layers make up only a small part of Jamba, which is a possible reason for smaller performance gains.
LinProj S6 GLUE DART CelebA SAMSum Spider
Avg. BLEU MET. Acc. R1 R2 RL Acc.
LoRA LoRA 65.5 52.9 73.0 88.5 56.4 33.5 47.9 90.7
SDT 67.7 53.1 73.0 88.4 56.5 33.5 48.0 89.8
Table 22: Performance comparison between SDT and LoRA on pretrained Jamba models. Bold numbers indicate the best performance for each task. We use Jamba-Tiny-319M to compare the performance of SDT and LoRA on GLUE (Wang et al., 2019), and CelebA (Liu et al., 2015) benchmarks. For all other datasets, we employ Jamba-Mini-52B. We report only the best setting out of three for each method.
Additional Results for LoRA+.
We extend our investigation to include LoRA+ (Hayou et al., 2024) with SDT and evaluate its performance against LoRA+ across various datasets on both Mamba-I and Mamba-II. The results, presented in Table 23, show that integrating SDT with LoRA+ enhances its effectiveness and achieves superior performance compared to using LoRA+ alone.
Model Mamba-I-130M Mamba-II-130M Mamba-II-1.3B
Dataset DART DART SAMSum Spider
Metric ($\uparrow$) METEOR BLEU METEOR BLEU R1 R2 RL Acc.
LoRA+ 70.06 50.91 69.78 49.14 49.83 26.09 41.66 73.75
LoRA+ & SDT 70.58 51.93 70.48 49.99 50.81 27.19 42.4 84.22
Table 23: Performance comparison between LoRA+ and SDT on Mamba-I and Mamba-II. Bold numbers indicate the best performance for each task. We test all experiments under various parameter settings (<0.4%) for both LoRA+ and SDT, and report the best values.